HUBBLE SPACE TELESCOPE DAILY REPORT #2556 PERIOD COVERED: 0000Z (UTC) 02/08/00 - 0000Z (UTC) 02/09/00 Daily Status Report as of 040/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD/MA1/MA2 8479 (Intrinsic Absorption Lines IN Seyfert 1 Galaxies) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to observe the origin of the intrinsic absorption lines in the UV spectra of Seyfert 1 galaxies and their connection to the emission-line and X-ray absorption. Low-resolution spectra of NGC 3227 will be obtained to look for low ionization absorption. The observations completed with no reported problems. 1.2 Completed Two Sets of WF/PC-2 8460 (Cycle 8 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations completed with no reported problems. 1.3 Completed WF/PC-2 8457 (UV Earth Flats) The WF/PC-2 was used to obtain a series of Ultraviolet Earth flats. The observations were completed with no reported problems. 1.4 Completed Two Sets of STIS/CCD 8411 (STIS CCD Spectroscopic Flats C-8) The Space Telescope Imaging Spectrograph (CCD) was used to obtain CCD flats in the spectroscopic mode. The observations were completed as scheduled, and no anomalies were reported. 1.5 Completed STIS/CCD 8202 (Bright Quasar Close Lensing Search) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of HCLQ1152+3 as part of a second generation survey to search for bright lensing quasars. With the restored PSF, new images will be considerably more sensitive to the presence of both close separation and faint lens components, and because STIS reaches two magnitudes fainter the PC, the new snapshot images will have much greater dynamic range. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed STIS/CCD 8439 (Bias Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no reported problems. 1.7 Completed WF/PC-2 8193 (Comparing the Hosts of High-Z Radio-Quiet Quasars to Lyman Break Galaxies) The WF/PC-2 was used to try to determine the underlying causes of the rapid cosmological evolution of the quasar population that is a mystery around which is woven many different strands of contemporary cosmology. The proposal completed nominally. 1.8 Completed STIS/CCD 8437 (Dark Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. 1.9 Completed Five Sets of STIS/CCD 8549 (POMS Test Proposal: Non-Scripted Parallel Proposal Continuation I) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude, non-scripted parallel observations as part of a POMS test proposal. The observations were completed as planned, and no anomalies were reported. 1.10 Completed WF/PC-2 8230 (Snapshot Survey of Dynamically Close Galaxy Pairs from z=0.1 to z=0.5) The WF/PC-2 was used to make snapshot survey observations of galactic cluster Gal-Clus-0. This is to obtain snapshot images of a large, well-defined, and dynamically selected sample of close galaxy pairs at moderate redshift. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed Two Sets of WF/PC-2 8194 (Evolution of High-Redshift Seyfert Galaxies) The WF/PC-2 was used to make observations of the high-redshifted Seyfert galaxies Gal-105259 and Gal-105200 in order to study their evolution. The observations will provide insights into the Seyfert phenomenon at intermediate redshift and will complement similar HST studies of normal galaxies, clusters and quasars. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed WF/PC-2 8451 (Cycle 8 Photometric Characterization) The WF/PC-2 was used to determine if any changes in the zeropoint, the spatial dependence of the zeropoint, or contamination rates have occurred, by comparing the data with baseline measurements for GRW+70D5824. The proposal completed with no anomalies. 1.13 Completed WF/PC-2 8216 (T Tauri Star Snapshot Survey II. Completing the CTTs Sample) The WF/PC-2 was used to make observations of the T-Tauri class star known as MHO-4 as part of Snapshot Survey II. This database will be of fundamental importance to understanding the nature of planet-forming environments and their frequency in the galaxy. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed STIS/CCD/MA1/MA2 8245 (Black Hole Accretion Outbursts in Soft X- Ray Transients) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of the optical and ultraviolet counterpart of a soft X-ray outburst region from a black hole accretion disk. A wavelength and a CCD flat calibration images were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed WF/PC-2 8436 (POMS Test Proposal: WFII Backup Parallel Archive Proposal II) The WF/PC-2 was used to make archive observations parallel to the prime science observations as part of a POMS test proposal. This is a POMS test proposal designed to simulate scientific plans. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 9 Successful: 9 Scheduled Re-acquisitions: 7 Successful: 7 2.2 FHST Updates: Scheduled: 18 Successful: 18 2.3 Operations Notes: Servicing Mission Observatory Verification (SMOV) activities continue to be monitored by the operations team from the Building 23 back-up facility, using CCS string "A". Per ROP SR-1A, the SSR EDAC error counter was reset once. An SI C&DH error was reset at 039/1315Z, using ROP NS-45. The STIS MCE-1 reset at 040/040354Z while the low voltage was on. The STIS flight software error counter was reset at 040/0506Z as directed by ROP NS-12. MAMA-1 will be recovered via nominal SMS commanding at approximately 040/1717Z when the high voltage will be turned on. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of SMOV and normal science observations and calibrations. /CAW